The Australian sheep dog is a breed renowned for its intelligence and work ethic. It is estimated that 35% of adult Australian sheep dogs weigh 65 pounds or more. A sample of 15 adult dogs is studied. What is the probability that exactly 5 of them weigh 65 lb or more?
Answer #1

Binomial Distribution

n = 15

p = 0.35

q = 1 - p = 0.65

P(X=5)=\binom{15}{5}\times 0.65^{10}\times 0.35^{5}=3003\times 0.0134627\times 0.005252=0.2123

So,

Answer is:

0.2123